Water Sport Inc. manufactures a small personal water tube used for children learning to swim. Management is now preparing detailed budgets for the third quarter, July through September, and has assembled the following information to assist:


a. The Marketing Department has estimated sales as follows for the remainder of the year (number of water tubes):

  July 7,300   October 3,700
  August 5,800   November 3,300
  September 4,800   December 2,800


The selling price of the water tubes is $60.


b. All sales are on account. Based on past experience, sales are expected to be collected in the following pattern:

  50 % in the month of sale
  45 % in the month following sale
  5 % uncollectible

The beginning accounts receivable balance (excluding uncollectible amounts) on July 1 will be $164,000.

  1. The company maintains finished goods inventories equal to 20% of the following month’s sales. The inventory of finished goods on July 1 will be 1,460 units.
  2. Each water tube requires 3 kilograms of synthetic polyisoprene rubber compound. To prevent shortages, the company would like the inventory of synthetic rubber compound on hand at the end of each month to be equal to 20% of the following month’s production needs. The inventory of synthetic rubber compound on hand on July 1 will be 4,200 kilograms.
  3. The synthetic rubber compound costs $3.00 per kilogram. Water Sport pays for 70% of its purchases in the month of purchase; the remainder is paid for in the following month. The accounts payable balance for synthetic rubber compound purchases will be $13,800 on July 1.
Answer #1
SALES AND CASH COLLECTION BUDGET
July August September October November December
A Sales in Units             7,300             5,800              4,800             3,700             3,300             2,800
B=A*$60 Sales Revenue $438,000 $348,000 $288,000 $222,000 $198,000 $168,000
C=B*50% Collection from the month of Sales $219,000 $174,000 $144,000 $111,000 $99,000 $84,000
D Collection from previous month of Sales(45%) $164,000 $197,100 $156,600 $129,600 $99,900 $89,100
E=C+D Total Cash Collection during the month $383,000 $371,100 $300,600 $240,600 $198,900 $173,100
Collection in August from July Sales = $197,100 45%*438000
Collection in July from Previous Sales = $164,000 Accounts Receivable
PRODUCTION BUDGET
July August September October November December
A Sales in Units             7,300             5,800              4,800             3,700             3,300             2,800
B End of Month inventory             1,160                 960                  740                 660                 560
C Beginning of Month Inventory             1,460             1,160                  960                 740                 660                 560
D=A+B-C Budgeted Production in units             7,000             5,600              4,580             3,620             3,200
End of Month inventory in July =20%*5800             1,160
Beginning of month inventory in July             1,460
DIRECT MATERIALS PURCHASE BUDGET
July August September October November
A Budgeted Production in units             7,000             5,600              4,580             3,620             3,200
B=A*3 Raw materials required (Kilograms)           21,000           16,800            13,740           10,860             9,600
C End of month inventory(Kilograms)             3,360             2,748              2,172             1,920
D Beginning of months inventory(Kilograms)             4,200             3,360              2,748             2,172
E=B+C-D Budgeted direct material purchase(Kilograms)           20,160           16,188            13,164           10,608
F=E*$3 Budgeted direct material purchase(Dollars) $60,480 $48,564 $39,492 $31,824
End of months inventory in July =20%*16800=             3,360
Beginning of months inventory in July             4,200
PURCHASE DISBURSEMENT BUDGET
July August September October
A Budgeted direct material purchase(Dollars) $60,480 $48,564 $39,492 $31,824
B=A*70% Payment for the month of purchase $42,336 $33,995 $27,644 $22,277
C Payment for previous months purchase $13,800 $18,144 $14,569 $11,848
D=B+C Budgeted Purchase Disbursement $56,136 $52,139 $42,214 $34,124
Payment in August for July Purchase=30%*60480 $18,144
Payment in July for Previous Purchase $13,800 Accounts Payable